# Periodicity of traces of Hecke operators modulo prime powers

### v1

@@ 1 Introduction
In this paper, we prove congruences, modulo prime powers, between traces of Hecke operators acting on spaces of classical cusp forms of different weights. Our result extends a result of Serre [serreletter], who determined such congruences modulo primes. The congruences hold for very general congruence subgroups; in particular, we do not assume that $p$-adic analytic methods apply, but nevertheless obtain results in the spirit of the Gouvêa–Mazur conjecture.
Our main tool is the trace formula for Hecke operators established in Deligne’s foundational work [Deligne]. Our methods extend naturally to the setting of Drinfeld modular forms over function fields, giving a parallel theory of Hecke trace congruences in positive characteristic.
@@ 1.1 The main theorems
If $H$ is not representable and $\ell=2$ or $3$, then the above statement holds but with $s$ replaced by $s+\nu_{\ell}(H,q)$ in the definition of $n$. In the case $\ell=2$ and $\mathrm{Id}\in H$ then $s$ can instead be replaced by $s-1+\nu_{2}(H,q)$ in the definition of $n$.
@@
In a letter to Tate (see Equation (2) and (5) of [serreletter], but note the differences in notation), Serre proved that if $w_{k}=\operatorname{Tr}(F_{p}\,|\,S[\Gamma,k])-\operatorname{Tr}(F_{p}\,|\,S[\Gamma,k-(\ell-1)])$, then $w_{k+\ell+1}\equiv_{\ell}p\cdot w_{k}$ for all $\Gamma=\Gamma(N)$ with $N\geq 3$, $q=p\neq\ell$, $s=1$ and $k>\ell+1$. Under these assumptions it follows that
@@
Theorem 1.1 in the case $\ell=p$ is related to the theory of $p$-adic modular forms. Let $p\geq 5$ and $p\nmid N$. Let $P_{k}(t)=\det(1-U(p)t\,|\,M_{k}(N;r))$ denote the characteristic power series of the Atkin $U(p)$-operator acting on $r$-overconvergent $p$-adic modular forms of weight $k$ and level $\Gamma_{1}(N)$, where $r\in\mathbb{Q}_{p}$ satisfies $0<\text{ord}_{p}(r)<p/(p+1)$. Gouvêa and Mazur showed in [GM, Theorem 1] that for any $s\geq 1$ and $k\in\mathbb{Z}$, one has the congruence $P_{k}(t)\equiv_{p^{s}}P_{k+p^{s-1}(p-1)}(t)$. Notice that the period is precisely the same as the one from Theorem 1.1, which shows that similar congruences hold when $k$ is large enough for $T(p)$ acting on spaces of classical modular forms of essentially arbitrary level. See also Proposition 5.5 for a statement about characteristic polynomials mod $p$.
@@ 2.1 Modular curves
Let now $H\subseteq\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ be a subgroup. Let $\mathcal{Y}_{H}$ (respectively $\mathcal{X}_{H}$) be the moduli space of elliptic curves (respectively generalized elliptic curves) with a level $H$-structure. This is a smooth (respectively smooth and proper) Deligne–Mumford stack defined over $\mathbb{Z}[1/N]$ with geometrically irreducible fibers over $\mathbb{Z}[\zeta_{N}]^{\det H}$, see [DeligneRapoport, Section IV.3 and IV.5]. In particular, $\mathcal{Y}_{H}(\mathbb{C})\cong\coprod_{i=1}^{d}\Gamma_{i}\backslash\mathcal{H}$, where $\mathcal{H}$ is the complex upper half plane and each $\Gamma_{i}$ is a congruence subgroup of $\mathrm{SL}_{2}(\mathbb{Z})$. Write $S_{k}(H)=\bigoplus_{i=1}^{d}S_{k}(\Gamma_{i})$ and $S[H,k]=\bigoplus_{i=1}^{d}S[\Gamma_{i},k]$. When $\det H=(\mathbb{Z}/N\mathbb{Z})^{\times}$, we have $S[H,k]=S[\Gamma,k]$ where $\Gamma=\pi^{-1}(H)$ for $\pi:\mathrm{SL}_{2}(\mathbb{Z})\to\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$. In particular, for $\Gamma\in\{\Gamma(N),\Gamma_{1}(N),\Gamma_{0}(N)\}$, one can find $H_{\Gamma}\subset\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ such that $S[H_{\Gamma},k]=S[\Gamma,k]$.
@@
If, for all fields $k$ with $\text{char}(k)\nmid N$ and all $y=(E,\varphi)\in\mathcal{Y}_{H}(k)$, the automorphism group $\operatorname{Aut}_{k}(y)$ is trivial, then we call $H$ representable. For instance, if $\pi^{-1}(H)\subset\Gamma(N)$, with $N\geq 3$, then $H$ is representable, see for instance [DeligneRapoport, Théorème IV.2.7].
@@
where $\nu_{\ell}$ is the $\ell$-adic valuation on $\mathbb{Z}$ normalised so that $\nu_{\ell}(\ell)=1$. Note that $\nu_{2}(H,q)\leq 3$, $\nu_{3}(H,q)\leq 1$ and $\nu_{\ell}(H,q)=0$ for $\ell\geq 5$, see [silverman, Theorem 10.1]. The elements of $(\mathcal{X}_{H}\setminus\mathcal{Y}_{H})(k)]$ are called cusps. Finally define the $\ell^{\prime}$-adic local system $\mathbb{V}_{k}=\mathrm{Sym}^{k}(R^{1}\psi_{*}\mathbb{Q}_{\ell^{\prime}})$ on $\mathcal{Y}_{H}$.
@@
$a_{1}^{k-2j}]=\sum_{[(E,\varphi)]\in\mathcal{[}Y_{H}(\mathbb{F}_{q})]}\frac{a_{1}(E)^{k-2j}}{\#\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)}\in\mathbb{Z},$
@@
Fix any $\ell^{\prime}\nmid q$. Below, we consider $\mathcal{X}_{H}$ over $\overline{\mathbb{F}}_{q}$ and we let $F_{q}$ denote the geometric Frobenius morphism. The Eichler–Shimura–Deligne relation together with the Grothendieck–Lefschetz trace formula, see [Deligne] and [Scholl, Proposition 4.4], says that for $k>0$
@@
Again using the Eichler–Shimura–Deligne relation and the Grothendieck-Lefschetz trace formula,
@@
Together with Lemma 2.4 we get that,
@@
If $\pi^{-1}(H)\subset\Gamma(M)$ with $M\geq 3$, then all cusps will be of split multiplicative type and so $\mathrm{eis}_{H,k}(q)$ will equal the number of $\overline{\mathbb{F}}_{q}$-isomorphism classes of $(\mathcal{X}_{H}\setminus\mathcal{Y}_{H})(\mathbb{F}_{q})$ for all $k$, see [Scholl, Proof of Proposition 4.4]. For $\pi^{-1}(H)=\mathrm{SL}_{2}(\mathbb{Z})$, we have one split cusp and one non-split cusp over $\mathbb{F}_{q}$, both with an $\mathbb{F}_{q}$-automorphism 1-1 of order $2$ (and the two cusps are twists by 1-1 of each other) giving $\mathrm{eis}_{H,k}(q)=(1^{k}+(-1)^{k})/2.$ For computations of $\mathrm{eis}_{H,k}(q)$ in the case when $\pi^{-1}(H)=\Gamma_{1}(3)$ or $\Gamma_{1}(4)$, see [Hoffmanetal, Theorem 2].
@@
For any $k\geq 1$ let $h_{k}(x_{1},x_{2})$ denote the complete homogeneous symmetric polynomial of weight $k$. Then
@@ 2.2 Drinfeld modular curves
Let $\mathfrak{p}$ be a non-zero prime ideal of $A[\mathfrak{n}^{-1}]$. Let $\mathbf{T}_{\mathfrak{p}}$ denote the $\mathfrak{p}$-Hecke operator acting on the space $\operatorname{S}_{k+2,l}({\mathcal{K}})$ of Drinfeld modular cusp forms of weight $k+2$, type $l$, and level $\mathcal{K}$}. There is then a decomposition
@@
The spaces $\operatorname{S}_{k+2,l}(\mathcal{K})$ admit actions of Hecke operators $\mathbf{T}^{\prime}_{\mathfrak{p}}$, see [bockle, Section 6]. For any $n\geq 1$, we have $\mathbf{T}^{\prime}_{\mathfrak{p}^{n}}=(\mathbf{T}^{\prime}_{\mathfrak{p}})^{n}$ (in contrast to Remark 2.3), and Böckle–Eichler–Shimura theory yields the equality [devries_ramanujan, Remark 4.12]
@@
where $\pi_{\varphi},\bar{\pi}_{\varphi}$ denote the roots of the Frobenius polynomial of $\varphi$ (see [papikian, Chapter 4.2] or [gek_finite]).
@@
The trace formula takes values in the characteristic $p$ ring $A$, so in particular only the residue classes $\operatorname{Aut}_{\mathbb{F}_{\mathfrak{p}^{n}}}(\varphi,\lambda)\pmod{p}$ play a role. The automorphism group of a Drinfeld module (without level structure) over $\mathbb{F}_{\mathfrak{p}^{n}}$ can be identified with the group of units in a finite field of the same cardinality, which has cardinality $1\pmod{p}$.
2.2.1 Renormalising the Hecke operators
Assume $\deg(\infty)=1$, so that the residue field $\mathbb{F}_{\infty}$ of $\infty$ is $\mathbb{F}_{q}$. Let $\pi_{\infty}$ be a uniformizer of the local ring $\mathcal{O}_{\infty}\subset K_{\infty}$ and let $\text{pr}:\mathcal{O}_{\infty}\to\mathbb{F}_{\infty}$ be the reduction modulo $\pi_{\infty}$. We say that $a\in K^{\times}$ is monic if $\text{pr}(\pi_{\infty}^{v_{\infty}(a)}a)=1$. By assumption, for any $a\neq 0$ there is a unique $b\in\mathbb{F}_{q}^{\times}$ such that $ba$ is monic.
Now assume $\mathfrak{M}_{\mathcal{K},\mathfrak{p}}(\mathbb{F}_{\mathfrak{p}^{n}})\neq\emptyset$. Then the ideal $\mathfrak{p}^{n}$ is principal [gek_finite, Cor. 5.2] and, by the above, has a unique monic generator, which we denote by $\wp_{n}$. In this case, $\wp_{nm}=\wp_{n}^{m}$ for any $m\geq 1$.
@@ 2.2.1 Renormalising the Hecke operators
    By definition of Drinfeld modular forms, if $l\equiv l^{\prime}\pmod{q-1}$ then $\operatorname{S}_{k,l}({\mathcal{K}})=\operatorname{S}_{k,l^{\prime}}({\mathcal{K}})$ as vector spaces. This identification is Hecke-equivariant with the new normalisation, cf. [bockle, Remarks 5.33 and 6.12].
@@
    If $\varphi\in\mathfrak{M}_{\mathcal{K},\mathfrak{p}}(\mathbb{F}_{\mathfrak{p}^{n}})$, then $\pi_{\varphi}\bar{\pi}_{\varphi}=b_{\varphi}\wp_{n}$ for a unique $b_{\varphi}\in\mathbb{F}_{q}^{\times}$. This simplifies the trace formula: (1) yields
@@
If $\mathfrak{p}^{n}$ is not a principal ideal, then the fact that there are no Drinfeld modules over $\mathbb{F}_{\mathfrak{p}^{n}}$ [gek_finite, Cor. 5.2] implies, by (2), that the trace of $\mathbf{T}_{\mathfrak{p}}^{n}$ is zero on any space of Drinfeld modular forms. Therefore we will always assume that $\mathfrak{p}^{n}$ is principal.
@@ 3 Proof of the main theorem for elliptic modular forms
    $\mathcal{N}(\mathbb{F}_{q}):=\{[(E,\varphi)]\in[\mathcal{Y}(\mathbb{F}_{q})]\,|\,a_{1}(E)\equiv_{\ell}0\},\qquad\mathcal{U}(\mathbb{F}_{q}):=[\mathcal{Y}(\mathbb{F}_{q})]\setminus\mathcal{N}(\mathbb{F}_{q}).$
@@
If $\ell\geq 5$ then $\ell\nmid\#\operatorname{Aut}_{\mathbb{F}_{q}}(E)$ for any elliptic curve $E$ over $\mathbb{F}_{q}$, see [silverman, Theorem 10.1]. For any level structure $\varphi$, $\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)$ is a subgroup of $\operatorname{Aut}_{\mathbb{F}_{q}}(E)$, which is trivial if $H$ is representable. Hence, under our assumptions on $\ell$ we have that $\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)$ is invertible mod $\ell^{s}$.
@@
This follows immediately from [prime-power-lucas, Theorem 1] [^1] In the statement of this theorem, $m=n+r$ should read $n=m+r$.. ∎
@@
$N_{j}(k+2):=\binom{(k+\delta_{k})/2+j}{2j+\delta_{k}}(-q)^{(k-\delta_{k})/2-j}\sum_{[(E,\varphi)]\in\mathcal{N}(\mathbb{F}_{q})}\frac{a_{1}(E)^{2j+\delta_{k}}}{\#\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)}$
By definition of $\mathcal{N}(\mathbb{F}_{q})$, we see that the sum over $\mathcal{N}(\mathbb{F}_{q})$ is divisible by $\ell^{k-2j}$, so it suffices to show that
@@
Since $\ell\nmid q$, we have that q-q is a unit mod $\ell^{s}$ and $(-q)^{n_{N}/2}\equiv_{\ell^{s}}1$ by definition of $n_{N}$. The required congruence of the binomial coefficients follows from Lemma 3.4. ∎

### v2

@@ 1 Introduction
In this paper, we prove congruences, modulo prime powers, between traces of Hecke operators acting on spaces of classical cusp forms of different weights. Our result extends a result of Serre [18], who determined such congruences modulo primes. The congruences hold for very general congruence subgroups; in particular, we do not assume that $p$-adic analytic methods apply, but nevertheless obtain results in the spirit of the Gouvêa–Mazur conjecture.
Our main tool is the trace formula for Hecke operators established in Deligne’s foundational work [7]. Our methods extend naturally to the setting of Drinfeld modular forms over function fields, giving a parallel theory of Hecke trace congruences in positive characteristic.
@@ 1.1 The main theorems
If $H$ is not representable and $\ell=2$ or $3$, then the above statement holds but with $s$ replaced by $s+\nu_{\ell}(H,q)$ in the definition of $n$. In the case $\ell=2$ and $\mathrm{Id}\in H$ then $s$ can instead be replaced by $s+\nu_{2}(H,q)-1$ in the definition of $n$.
@@
In a letter to Tate (see Equation (2) and (5) of [18], but note the differences in notation), Serre proved that if $w_{k}=\operatorname{Tr}(F_{p}\,|\,S[\Gamma,k])-\operatorname{Tr}(F_{p}\,|\,S[\Gamma,k-(\ell-1)])$, then $w_{k+\ell+1}\equiv_{\ell}p\cdot w_{k}$ for all $\Gamma=\Gamma(N)$ with $N\geq 3$, $q=p\neq\ell$, $s=1$ and $k>\ell+1$. Under these assumptions it follows that
@@
Let $p\geq 5$. Theorem 1.1 in the case $\ell=p$ was previously proved for level $\mathrm{SL}_{2}(\mathbb{Z})$ in [14, 1]. At level $\Gamma_{1}(N)$ with $p\nmid N$, it is related to the theory of $p$-adic modular forms. Let $P_{k}(t)=\det(1-U(p)t\,|\,M_{k}(N;r))$ denote the characteristic power series of the Atkin $U(p)$-operator acting on $r$-overconvergent $p$-adic modular forms of weight $k$ and level $\Gamma_{1}(N)$, where $r\in\mathbb{Q}_{p}$ satisfies $0<\text{ord}_{p}(r)<p/(p+1)$. Gouvêa and Mazur showed in [9, Theorem 1] that for any $s\geq 1$ and $k\in\mathbb{Z}$, one has the congruence $P_{k}(t)\equiv_{p^{s}}P_{k+p^{s-1}(p-1)}(t)$. Theorem 1.1 gives an analogous result for $T(p)$ acting on spaces of classical cusp forms of essentially arbitrary level, when $k$ is large enough. In the case of $\Gamma_{1}(N)$ this can be seen as a statement about $U(p)$ restricted to the $p$-old classical cusp forms.
See also Proposition 5.5 for a statement about characteristic polynomials mod $p$.
@@ 2.1 Modular curves
Let now $H\subseteq\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ be a subgroup. Let $\mathcal{Y}_{H}$ (respectively $\mathcal{X}_{H}$) be the moduli space of elliptic curves (respectively generalized elliptic curves) with a level $H$-structure. This is a smooth (respectively smooth and proper) Deligne–Mumford stack defined over $\mathbb{Z}[1/N]$ with geometrically irreducible fibers over $\mathbb{Z}[\zeta_{N}]^{\det H}$, see [6, Section IV.3 and IV.5]. In particular, $\mathcal{Y}_{H}(\mathbb{C})\cong\coprod_{i=1}^{d}\Gamma_{i}\backslash\mathcal{H}$, where $\mathcal{H}$ is the complex upper half plane and each $\Gamma_{i}$ is a congruence subgroup of $\mathrm{SL}_{2}(\mathbb{Z})$. Write $S_{k}(H)=\bigoplus_{i=1}^{d}S_{k}(\Gamma_{i})$ and $S[H,k]=\bigoplus_{i=1}^{d}S[\Gamma_{i},k]$. When $\det H=(\mathbb{Z}/N\mathbb{Z})^{\times}$, we have $S[H,k]=S[\Gamma,k]$ where $\Gamma=\pi^{-1}(H)$ for $\pi:\mathrm{SL}_{2}(\mathbb{Z})\to\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$. In particular, for $\Gamma\in\{\Gamma(N),\Gamma_{1}(N),\Gamma_{0}(N)\}$, one can find $H_{\Gamma}\subset\mathrm{GL}_{2}(\mathbb{Z}/N\mathbb{Z})$ such that $S[H_{\Gamma},k]=S[\Gamma,k]$.
@@
If, for all fields $k$ with $\text{char}(k)\nmid N$ and all $y=(E,\varphi)\in\mathcal{Y}_{H}(k)$, the automorphism group $\operatorname{Aut}_{k}(y)$ is trivial, then we call $H$ representable. For instance, if $H\subseteq H_{\Gamma(N)}$ with $N\geq 3$ then $H$ is representable, and if $6\mid N$ then $H$ is representable if and only if $\pi^{-1}(H)$ is torsion-free [6, Théorèmes IV.2.7, VI.2.7].
@@
where $\nu_{\ell}$ is the $\ell$-adic valuation on $\mathbb{Z}$ normalized so that $\nu_{\ell}(\ell)=1$. Note that $\nu_{2}(H,q)\leq 3$, $\nu_{3}(H,q)\leq 1$, and $\nu_{\ell}(H,q)=0$ for $\ell\geq 5$, see [19, Theorem 10.1]. The elements of $(\mathcal{X}_{H}\setminus\mathcal{Y}_{H})(k)]$ are called cusps. Finally define the $\ell^{\prime}$-adic local system $\mathbb{V}_{k}=\mathrm{Sym}^{k}(R^{1}\psi_{*}\mathbb{Q}_{\ell^{\prime}})$ on $\mathcal{Y}_{H}$.
@@
$a_{1}^{k-2j}]=\sum_{[(E,\varphi)]\in\mathcal{[}\mathcal{Y}_{H}(\mathbb{F}_{q})]}\frac{a_{1}(E)^{k-2j}}{\#\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)}\in\mathbb{Z},$
@@
Fix any $\ell^{\prime}\nmid q$. Below, we consider $\mathcal{X}_{H}$ over $\overline{\mathbb{F}}_{q}$ and we let $F_{q}$ denote the geometric Frobenius morphism. The Eichler–Shimura–Deligne relation together with the Grothendieck–Lefschetz trace formula, see [7] and [16, Proposition 4.4], says that for $k>0$,
@@
Again using the Eichler–Shimura–Deligne relation and the Grothendieck–Lefschetz trace formula,
@@
Together with Lemma 2.4 we get that
@@
For $N=1$, there is one split cusp and one non-split cusp over $\mathbb{F}_{q}$, both with an $\mathbb{F}_{q}$-automorphism 1-1 of order $2$ (and the two cusps are twists by 1-1 of each other) giving $\mathrm{eis}_{H,k}(q)=(1^{k}+(-1)^{k})/2.$ For computations of $\mathrm{eis}_{H,k}(q)$ in the case when $\pi^{-1}(H)=\Gamma_{1}(3)$ or $\Gamma_{1}(4)$, see [11, Theorem 2].
@@
For any $k\geq 1$, let $h_{k}(x_{1},x_{2})$ denote the complete homogeneous symmetric polynomial of weight $k$. Then
@@ 2.2 Drinfeld modular curves
Let $\operatorname{S}_{k+2,l}({\mathcal{K}})$ denote the space of Drinfeld modular cusp forms of weight $k+2$, type $l$, and level $\mathcal{K}$}. There is then a decomposition
@@
The spaces $\operatorname{S}_{k+2,l}(\mathcal{K})$ admit actions of Hecke operators $\mathbf{T}^{\prime}_{\mathfrak{p}}$, see [2, Section 6]. For any $n\geq 1$, we have $\mathbf{T}^{\prime}_{\mathfrak{p}^{n}}=(\mathbf{T}^{\prime}_{\mathfrak{p}})^{n}$ (in contrast to Remark 2.3), and Böckle–Eichler–Shimura theory yields the equality [4, Remark 4.12]
@@
where $\pi_{\varphi},\bar{\pi}_{\varphi}$ denote the roots of the Frobenius polynomial of $\varphi$ (see [15, Chapter 4.2] or [8]).
@@
The trace formula takes values in the characteristic $p$ ring $A$, so in particular only the residue classes $\operatorname{Aut}_{\mathbb{F}_{\mathfrak{p}^{n}}}(\varphi,\lambda)\pmod{p}$ play a role. The automorphism group of a Drinfeld module (without level structure) over $\mathbb{F}_{\mathfrak{p}^{n}}$ can be identified with the group of units in a finite field of the same characteristic, which has cardinality $1\pmod{p}$.
2.2.1 Renormalizing the Hecke operators
Assume from now on that $\deg(\infty)=1$, so that the residue field $\mathbb{F}_{\infty}$ of $\infty$ is $\mathbb{F}_{q}$. Fix a uniformizer $\pi_{\infty}$ of the local ring $\mathcal{O}_{\infty}\subset K_{\infty}$ and let $\text{pr}:\mathcal{O}_{\infty}\to\mathbb{F}_{\infty}$ be the reduction modulo $\pi_{\infty}$. We then say that $a\in K^{\times}$ is monic if $\text{pr}(\pi_{\infty}^{-v_{\infty}(a)}a)=1$. By assumption, for any $a\neq 0$ there is a unique $b\in\mathbb{F}_{q}^{\times}$ such that $ba$ is monic.
Now assume $\mathfrak{M}_{\mathcal{K},\mathfrak{p}}(\mathbb{F}_{\mathfrak{p}^{n}})\neq\emptyset$. Then the ideal $\mathfrak{p}^{n}$ is principal [8, Cor. 5.2] and, by the above, has a unique monic generator, which we denote by $\wp_{n}$. In this case, $\wp_{nm}=\wp_{n}^{m}$ for any $m\geq 1$.
@@ 2.2.1 Renormalizing the Hecke operators
    By definition of Drinfeld modular forms, if $l\equiv l^{\prime}\pmod{q-1}$ then $\operatorname{S}_{k,l}({\mathcal{K}})=\operatorname{S}_{k,l^{\prime}}({\mathcal{K}})$ as vector spaces. This identification is Hecke-equivariant with the new normalization, cf. [2, Remarks 5.33 and 6.12].
@@
    If $\varphi\in\mathfrak{M}_{\mathcal{K},\mathfrak{p}}(\mathbb{F}_{\mathfrak{p}^{n}})$, then $\pi_{\varphi}\bar{\pi}_{\varphi}=b_{\varphi}\wp_{n}$ for a unique $b_{\varphi}\in\mathbb{F}_{q}^{\times}$. Thus (1) yields
@@
If $\mathfrak{p}^{n}$ is not a principal ideal, then the fact that there are no Drinfeld modules over $\mathbb{F}_{\mathfrak{p}^{n}}$ [8, Cor. 5.2] implies, by (2), that the trace of $\mathbf{T}_{\mathfrak{p}}^{n}$ is zero on any space of Drinfeld modular forms. Therefore we will always assume that $\mathfrak{p}^{n}$ is principal.
@@ 3 Proof of the main theorem for elliptic modular forms
    $\mathcal{N}(\mathbb{F}_{q}):=\{[(E,\varphi)]\in[\mathcal{Y}_{H}(\mathbb{F}_{q})]\,|\,a_{1}(E)\equiv_{\ell}0\},\qquad\mathcal{U}(\mathbb{F}_{q}):=[\mathcal{Y}_{H}(\mathbb{F}_{q})]\setminus\mathcal{N}(\mathbb{F}_{q}).$
@@
If $\ell\geq 5$ then $\ell\nmid\#\operatorname{Aut}_{\mathbb{F}_{q}}(E)$ for any elliptic curve $E$ over $\mathbb{F}_{q}$, see [19, Theorem 10.1]. For any level structure $\varphi$, $\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)$ is a subgroup of $\operatorname{Aut}_{\mathbb{F}_{q}}(E)$, which is trivial if $H$ is representable. Hence, under our assumptions on $\ell$ we have that $\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)$ is invertible mod $\ell^{s}$.
@@
This follows immediately from [10, Theorem 1] [^1] In the statement of this theorem, $m=n+r$ should read $n=m+r$.. ∎
@@
$N_{j}(k+2):=\binom{(k+\delta_{k})/2+j}{2j+\delta_{k}}(-q)^{(k-\delta_{k})/2-j}\sum_{[(E,\varphi)]\in\mathcal{N}(\mathbb{F}_{q})}\frac{a_{1}(E)^{2j+\delta_{k}}}{\#\operatorname{Aut}_{\mathbb{F}_{q}}(E,\varphi)}.$
By definition of $\mathcal{N}(\mathbb{F}_{q})$, we see that the sum over $\mathcal{N}(\mathbb{F}_{q})$ is divisible by $\ell^{2j+\delta_{k}}$, so it suffices to show that
@@
Since $\ell\nmid q$, we have that q-q is a unit mod $\ell^{s}$ and $(-q)^{n_{N}/2}\equiv_{\ell^{s}}1$ by definition of $n_{N}$. The required congruence between the binomial coefficients follows from Lemma 3.4. ∎